# Unfolding the kagome lattice to improve understanding of ARPES in CoSn

### v1

@@ Abstract
Metallic kagome lattices are attracting significant attention as they provide a platform to explore the interplay between topology and magnetism. Angle-resolved photoemission spectroscopy (ARPES) plays a key role in unraveling their electronic structure. However, the analysis is often challenging due to the presence of multiple bands near the Fermi level. Indeed, each orbital generates three bands in a kagome lattice due to its three inequivalent sites, which soon becomes complicated if many orbitals are present. To address this complexity, using ARPES matrix elements can be highly beneficial. First, band symmetry can be determined through selection rules based on light polarization. We emphasize that, in kagome lattices, symmetry is not only determined by the orbital character but also by the relative phase between the three sublattices. Additionally, interference between the three sublattices leads to a strong modulation of ARPES intensity across neighboring Brillouin zones. We show how unfolded band calculations capture these modulations, helping with band identification. We apply these ideas to CoSn, whose simple structure retains the key features of a kagome lattice. Taking advantage of these two effects, we isolate the dispersion of each band and discuss novel correlation effects, selectively renormalizing the bands crossing the Fermi level and shifting the others.
@@ I Introduction
Metallic compounds with kagome lattices offer an exceptional means to study the interplay between a topological band structure, inherent to the kagome lattice Guo and Franz (2009); Mazin et al. (2014), and magnetism, ubiquitous in 3d kagome systems Yin et al. (2022); Wang et al. (2024). Interesting properties of magnetic kagome materials include a large anomalous Hall effect in the magnetic Weyl semimetal Co3Sn2S2 Liu et al. (2018), a Chern quantum phase in TbMn6Sn6 Yin et al. (2020), and possibly large pseudomagnetic fields in strained FeSn Zhang et al. (2023). Other electronic instabilities, such as exotic charge density waves in the non-magnetic AV3Sb5 Guo et al. (2024) and magnetic FeGe Teng et al. (2022), are also under intense investigation. ARPES plays a crucial role in characterizing the electronic structure of these materials Ye et al. (2018); Kang et al. (2020a, b); Lin et al. (2020); Liu et al. (2020); Hu et al. (2023); Peng et al. (2024), allowing visualization of Dirac cones, van Hove singularities, and flat bands, characteristic of the kagome band structure. Moreover, it is well suited to characterize the strength of electronic correlations, appearing, for example, as renormalization of the calculated band dispersion. Most metallic kagomes are found to be weakly correlated, with renormalization factors less than 2, but some exhibit more correlations, like CsCr3Sb5 Peng et al. (2024). In these intermetallic kagome systems, a problem is often that the weak crystal field does not significantly lift the degeneracy of the five orbitals. This leads to a rather complex situation with 15 bands often strongly overlapping near the Fermi level (5 per orbital times the 3 kagome sites).
@@
Fortunately, there are some ways to separate bands in an ARPES experiment. The most common method is to use selection rules associated with linear polarization to select orbitals that are either odd or even with respect to mirror planes containing high symmetry directions Damascelli et al. (2003); Moser (2023). In single-site systems, this depends solely on the orbital parity with respect to that mirror plane. In multi-site systems (e.g., non-Bravais lattices), such as the kagome lattice, this also depends on the relative phases of the orbitals at different sites Brouet et al. (2012), which is not always properly taken into account in the ARPES literature. We will detail these selection rules for a kagome lattice.
@@
Another potentially useful selection rule is the intensity modulation of the bands in different Brillouin Zones (BZ), due to interference between the atoms of the motif. This effect is similar to the form factor in structural experiments, which causes extinctions of some Bragg peaks in non-Bravais lattices. It naturally appears in ARPES matrix elements Nishimoto et al. (1996); Moser (2023), but it is much less utilized. For a 2-site motif, there will typically be one bonding and one antibonding band. The interference is constructive only for the bonding band, yielding apparently different dispersion in the first and second BZ. This can, for example, be observed in perovskites with rotated octahedra, like iridates Martins et al. (2018) or ruthenates Sohn and Kim (2022), but also with some more subtleties in graphene Liu et al. (2011) or iron pnictides Brouet et al. (2012). The case of the 3 sites lattice has been much less studied, and it is less predictable. We show here that unfolded calculations Dirnberger et al. (2021); Rubel et al. (2014) can easily predict these effects in complex systems. The relevance of unfolded calculations to analyze ARPES was already indicated for supercell systems Ku et al. (2010). More surprisingly, it can also be applied to systems like graphene Lee et al. (2018), where the definition of supercell is less intuitive. We show that this can be rationalized within an extended zone scheme. We use here unfolded calculations to obtain the spectral weight of the different bands in different BZ. Although it is not the only contribution to ARPES matrix elements, it often has a significant impact.
@@
We illustrate the benefit of taking these two points (polarization and the use of different BZ) into account with the example of CoSn, a relatively simple metallic and non-magnetic kagome system. Its electronic structure was already mapped out by ARPES Liu et al. (2020); Kang et al. (2020a); Cheng et al. (2023), but we pay particular attention to the role of matrix elements. By employing different polarizations along high-symmetry directions, we achieve a complete separation between odd and even bands. We further observe strong modulation of the bands’ intensities in different BZs, well predicted by unfolded calculations. We finally discuss the impact of correlations on each band and suggest different behaviors for the bands crossing the Fermi level and the filled ones at higher binding energies. This methodology should be increasingly helpful for more complex kagome systems that deviate more strongly from DFT calculations.
@@ II Methods
Single crystals were grown from a Sn-flux method, as in reference Liu et al., 2020. Their structure was characterized by X-ray diffraction, their composition by Energy Dispersion X-ray analysis and their electronic properties by transport measurements. ARPES experiments were carried out on UHV-cleaved single crystals at the CASSIOPEE beamline of SOLEIL synchrotron and the BLOCH beamline of MAX-IV synchrotron, at low temperatures (T15 K) and with an overall resolution better than 15meV.
DFT calculation were done using the full-potential linear augmented plane wave method implemented in the WIEN2k software Blaha et al. (2001). The experimental crystal structure (space group P6/mmm, No. 191) of CoSn with lattice constants Å and Å was used in the calculations.
@@ III.1 Tight-binding model
Let us label the central atom as 1 and the two orthogonal axes as and . The two planes and , which are perpendicular to the kagome plane, are mirror planes of the structure. They correspond respectively to K and M directions of the BZ (see inset in Fig. 1). One can see that reflection with respect to these mirror planes exchanges atoms 2 and 3 while leaving atom 1 unchanged. Therefore, the three bands created from a single orbital must combine atoms 2 and 3 in a symmetric () or antisymmetric () way to respect the mirror planes. In the first case, a contribution from atom 1 can be added (in-phase or out-of-phase), while in the second case, it must be zero.
We take as an example a orbital at site 1, which is even with respect to the mirror plane and odd with respect to the mirror plane. To preserve the symmetry of the kagome lattice, the orbital must rotate by 120∘ at sites 2 and 3. This is achieved by taking an appropriate combination of and orbitals Kim and Liu (2023). In other words, we consider the even component of the / combination. The typical dispersion of the three bands obtained from a tight-binding (TB) model with only first-neighbor couplings Guo and Franz (2009) is shown in Fig. 1(b). We neglect here spin-orbit coupling (SOC) that will open a gap where two bands touch at or K Guo and Franz (2009). This is crucial to analyse topological properties, but will not play a large role away from the crossing, where we will focus to identify the bands character. We further indicate on the figure the relative contributions of the three atoms in the first Brillouin zone, and we use color to represent the band parity with respect to mirror planes. The parity matches that of the orbital when sites 2 and 3 are in phase, and it is opposite when they are out-of-phase. Notably, forming a Dirac cone at K requires bands of opposite parity. However, it is often incorrectly assumed in ARPES studies of kagome compounds that band parity can be inferred directly from the dominant orbital character (leading to a wrong application of the selection rules that we describe below).
@@ III.2 CoSn calculated band structure
CoSn has a relatively simple structure with kagome planes stacked vertically on top of each other and separated by two Sn2 buffer layers Sales et al. (2021). This preserves the characteristic features of the simple TB model of the kagome plane, especially Dirac cones and relatively flat bands, as can be seen in the DFT calculation Liu et al. (2020) at of Fig. 2(a). More precisely, each orbital displays a 3-band structure with very well-defined orbital character (see symbols). We label 15 the bands built from / orbitals and 710 bands from /, observable in the chosen energy window. The orbitals appear only below eV at =0 and will be ignored for the rest of the discussion (at =, a strong three-dimensional dispersion brings it much closer to the Fermi level). Two Dirac cones are formed around eV for / (bands 3 and 4) and / (bands 8 and 9). Two nearly flat bands are present near eV at for (band 7) and (band 2), with a dispersion of about 0.2 eV, much smaller than the 1 eV dispersion of the Dirac bands.
The three bands are clearly visible in this energy window, as highlighted by the marker size in Fig. 2(b). Some residual weight is observed along bands 1 and 2, which are predominantly from , evidencing hybridization between the two. The symmetry along M and K is described by the point group. This symmetry is characterized by two perpendicular mirror planes: the kagome plane itself and the perpendicular plane containing the considered symmetry direction. Since the symmetry with respect to this latter mirror plane defines the ARPES selection rules Damascelli et al. (2003), we group in Fig. 2(b) the four irreducible representations of into two groups: those that are even (red) and those that are odd (blue) with respect to this mirror plane. One can check that they follow the same rules as in the TB model of Fig. 1(b).
@@ III.3 CoSn measured band structure with ARPES
We now turn to the comparison with actual ARPES experiments. In Fig. 3(a), we show a Fermi Surface for a CoSn sample lightly doped with Fe (5%), taken with linear horizontal (LH) polarization and at 126 eV photon energy, corresponding to =0 Kang et al. (2020a). In ARPES, polarization selection rules apply when the photoelectron momentum lies in high-symmetry planes, in this case and containing either K or M depending on the sample orientation. If the polarization is even (odd) with respect to this plane, only even (odd) bands will have non-zero matrix elements Damascelli et al. (2003); Moser (2023). Fig. 3(b) sketches the experimental geometry with the orientation of LH and LV polarization in our experiment. As LH lies within the plane, it is even with respect to it, and even bands will be detected along . As LH has both even and odd components with respect to , no straightforward selection rules can be applied. On the other hand, LV polarization is along and perpendicular to , selecting even bands along and odd ones along . Note that these selection rules only predict if a band can be observed and not if its intensity will be strong.
@@ IV Unfolding the kagome lattice
This will directly impact ARPES measurements in different BZs, as it was shown for other multisite systems like graphite or graphene Nishimoto et al. (1996); Moser (2017) that the intensity of the ARPES bands is modulated by the sum of the TB coefficients (see also the Appendix C).
Within the TB framework, this sum can easily be calculated and is represented by marker size in Fig. 4(b-d), for each inequivalent path. One can see that only one band is strong for each of the paths, specifically the one with the three atoms in phase (see also Appendix A, which describes the same effect as a function of the hidden periodicities). This emphasizes the need to measure ARPES at least along these three paths to observe all bands. In contrast, this is a means to better identify the nature of the band observed.
More generally, when a tight-binding (TB) model is either intractable or fails to capture the relevant physics, the electronic structure can be computed at the DFT level. The resulting wave functions can then be “unfolded” Dirnberger et al. (2021); Popescu_PRB_85_2012; Rubel et al. (2014) by projecting the states obtained in the true unit cell onto the reciprocal space of an alternative unit cell that accounts for additional periodicities
Such procedures are now commonly used to compare theoretical calculations with ARPES measurements when a supercell arises from an electronic ordering, such as antiferromagnetism or charge density waves. The resulting folded bands, with periodicity imposed by electronic order, typically exhibit weak intensity, in fact proportional to the strength of the perturbation Voit et al. (2000); Brouet et al. (2004). This approach can also be beneficial in calculations that involve large unit cells, for example, to simulate doping effects Berlijn et al. (2012).
We propose here to use it to analyze the kagome lattice, where the origin of the subperiodicities is intrinsic to the non-Bravais lattice type (see Appendix A). To unfold the band structure in the large BZ (dotted line in Fig. 4(a)), we use the Fold2Bloch module Rubel et al. (2014, 2023) and a transformation matrix . Figure 5 displays this calculation unfolded along the three different paths identified in Fig. 4(a). The size of the marker gives the unfolded weight corresponding to Eq. 2 and the colors refer to the orbital character, as in Fig. 2. There are well defined differences expected for spectral weights in different BZ, which will be the largest when all atoms are in-phase. For example band 1 is very strong in (a), while bands 3 and 4 are nearly zero. The opposite is true in panel (b) and (c), but differently for M and K. Generally, the unfolding effect is less clear on / orbitals, because they are more strongly hybridized with each other, blurring the differences in atomic composition ( and having opposite parities, they can only be hybridized for different atomic combinations).
@@ V Electronic correlations in CoSn
Figure 6(a) gives an overview of almost all the bands, by adding images measured in different experimental conditions in pure CoSn. Five bands are close together at around 0.4 eV and it is useful to use the combination of unfolding and parity outlined in Fig. 3 to separate them. Although SOC mixes the bands in the crossing region, blurring there the definition of parity and orbital character, they remain well enough defined away from to perform this analysis. On the other hand, this mixing is at the origin of the non-trivial topology of CoSn, which was precisely studied in ref. Kang et al., 2025. A zoom on the 0.6 eV region below is shown in Fig. 6(b,c) with models of the experimental dispersion as symbols. Our findings are in very good agreement with previous ARPES studies of CoSn Liu et al. (2020); Kang et al. (2020a); Cheng et al. (2023), which all find a bottom position of band 1 near 0.2 eV, the flat band 7 near 0.05 eV at K and the Dirac cones around 0.8 eV. Qualitatively, this agrees well with the calculated band structure, including SOC, which is shown as thin lines in Fig. 6(a). Quantitatively, the calculated dispersions are in fact shifted up by 0.18 eV, evidencing some deviation. The calculated bottom of band 1 at 0.37 eV is significantly different from the experimental value.
In principle, one would expect electronic correlations to induce a renormalization of the dispersions rather than an energy shift, as the latter does not conserve the number of carriers. Indeed, the Fermi level crossing of band 1 is not well described in Fig. 6(a) (the green arrows indicate the deviation). In Fig. 6(b), band 1 is renormalized by a factor of 2, which describes very well both the bottom of the band and its . This indicates that our data agree with the expected CoSn stoichiometry at . However, if all bands are renormalized by a factor of 2, as done in Fig. 6(b), the agreement worsens at higher binding energies. Band 3 is 0.05 eV lower in experiment than in this fit, and it is more dispersive [see green arrow in Fig. 6(b)]. At even higher binding energies [see Fig. 6(a)], band 4 clearly disperses from 1.35 eV at to 0.8 eV at K, which cannot be described by the renormalized band structure. We observe that the splitting between band 1 and 3 is 0.1 eV, exactly as in the unrenormalized band dispersion with SOC. This suggests that only the band crossing the Fermi level is renormalized. As it is tied by symmetry to other bands (band 1 and band 2 are only separated by the SOC gap, for example), renormalizing band 1 necessarily raises band 2 and the bands strongly hybridized with it. In Fig. 6(c), we renormalize band 1 by 2 and shift up all other bands by 0.18 eV, which gives a reasonable fit, although further orbital-dependent adjustments may be necessary. We note that a rather large range of renormalization values was given for CoSn, between 1.2 Kang et al. (2020a); Cheng et al. (2023) and 2 Liu et al. (2020), probably depending on which features were predominantly fitted.

### v2

@@ Abstract
Metallic kagome lattices are attracting significant attention as they provide a platform to explore the interplay between topology and magnetism. Angle-resolved photoemission spectroscopy (ARPES) plays a key role in unraveling their electronic structure. However, the analysis is often challenging due to the presence of multiple bands near the Fermi level. Indeed, each orbital generates three bands in a kagome lattice due to its three sites motif, which soon becomes complicated if many orbitals are present. To address this complexity, using ARPES matrix elements can be highly beneficial. First, band symmetry can be determined through selection rules based on light polarization. We emphasize that, in kagome lattices, as in all multi-site lattices, symmetry of the Bloch state is not only determined by the orbital character but also by the relative phase between the three sublattices. Additionally, interference between the three sublattices leads to a strong modulation of ARPES intensity across neighboring Brillouin zones. We show how unfolded band calculations capture these modulations, helping with band identification. We apply these ideas to CoSn, whose simple structure retains the key features of a kagome lattice. Using polarization dependent ARPES in several Brillouin zones, we isolate the dispersion of each band and discuss novel correlation effects, selectively renormalizing the bands crossing the Fermi level and shifting the others.
@@ I Introduction
Metallic compounds with kagome lattices offer an exceptional means to study the interplay between a topological band structure, inherent to the kagome lattice [1, 2], and magnetism, frequent in 3d kagome systems [3, 4]. Interesting properties of magnetic kagome materials include a large anomalous Hall effect in the magnetic Weyl semimetal Co3Sn2S2 [5], a Chern quantum phase in TbMn6Sn6 [6], and possibly large pseudomagnetic fields in strained FeSn [7]. Other electronic instabilities, such as exotic charge density waves in the non-magnetic AV3Sb5 [8] and magnetic FeGe [9], are also under intense investigation. ARPES plays a crucial role in characterizing the electronic structure of these materials [10, 11, 12, 13, 14, 15, 16], allowing location of Dirac cones, van Hove singularities, and flat bands, characteristic of the kagome band structure. Moreover, it is well suited to characterize the strength of electronic correlations, appearing, for example, as renormalization of the calculated band dispersion. Most metallic kagomes are found to be weakly correlated, with renormalization factors less than 2, but some exhibit more correlations, like CsCr3Sb5 [16]. In these intermetallic kagome systems, a problem is often that the weak crystal field does not significantly lift the degeneracy of the five orbitals. This leads to a rather complex situation with 15 bands often strongly overlapping near the Fermi level (5 per orbital times the 3 kagome sites).
@@
Fortunately, there are some ways to separate bands in an ARPES experiment. The most common method is to use selection rules associated with linear polarization to select orbitals that are either odd or even with respect to mirror planes containing high symmetry directions [17, 18]. In single-site systems, this depends solely on the orbital parity with respect to that mirror plane. In multi-site systems, such as the kagome lattice, this also depends on the relative phases of the orbitals at different sites [19], which is not always properly taken into account in the ARPES literature. We will detail these selection rules for a kagome lattice.
@@
Another potentially useful selection rule is the intensity modulation of the bands in different Brillouin Zones (BZ), due to interference between the atoms of the motif. This effect is similar to the form factor in structural experiments, which causes extinctions of some Bragg peaks in multi-site lattices. It naturally appears in ARPES matrix elements [20, 18], but it is much less utilized. For a 2-site motif, there will typically be one bonding and one antibonding band. The interference is constructive only for the bonding band, yielding apparently different dispersion in the first and second BZ. This was first observed in graphite [21], but also in perovskites with rotated octahedra, like iridates [22] or ruthenates [23] and, with some more subtleties, in graphene [25, 26, 24] or iron pnictides [19]. The case of the 3 sites lattice has been much less studied, and it is less predictable. We show here that unfolded calculations [27, 28] can easily predict these effects in complex systems. The relevance of unfolded calculations to analyze ARPES was already indicated for supercell systems [29]. More surprisingly, it can also be applied to systems like graphene [30], where the definition of supercell is less intuitive. We show that this can be rationalized within an extended zone scheme. We use here unfolded calculations to obtain the spectral weight of the different bands in different BZ. Although it is not the only contribution to ARPES matrix elements, it often has a significant impact.
@@
We illustrate the benefit of taking these two points (polarization and the use of different BZ) into account with the example of CoSn, a relatively simple metallic and non-magnetic kagome system. Its electronic structure was already mapped out by ARPES [14, 11, 31], but we pay particular attention to the role of matrix elements. By employing different polarizations along high-symmetry directions, we achieve a complete separation between odd and even bands. We further observe strong modulation of the bands’ intensities in different BZs, well predicted by unfolded calculations. We finally discuss the impact of correlations on each band and suggest different behaviors for the bands crossing the Fermi level and the filled ones at higher binding energies. This methodology should be increasingly helpful for more complex kagome systems that deviate more strongly from DFT calculations.
@@ II Methods
Single crystals were grown from a Sn-flux method, as in reference 14. Their structure was characterized by X-ray diffraction, their composition by Energy Dispersion X-ray analysis and their electronic properties by transport measurements [32]. ARPES experiments were carried out on UHV-cleaved single crystals at the CASSIOPEE beamline of SOLEIL synchrotron and the BLOCH beamline of MAX-IV synchrotron, at low temperatures (T15 K) and with an overall resolution better than 15meV.
DFT calculations were performed using the full-potential linearized augmented plane-wave method as implemented in WIEN2k [33], together with the generalized-gradient approximation for the exchange-correlation functional of Perdew, Burke, and Ernzerhof [34]. The plane-wave expansion in the interstitial region was determined by setting the product (the smallest muffin-tin radius and the largest wave vector). The Brillouin zone was sampled using a -centred Monkhorst-Pack -mesh, and the tetrahedron method was employed for the integration. Self-consistency convergence criteria were set to Ry for total energy and e for charge. Calculations were carried out using the experimental crystal structure of CoSn (space group , No. 191) with lattice parameters Å and Å. The muffin-tin radius was set to 2.49 bohr for Co and Sn.
@@ III.1 Tight-binding model
Let us label the central atom as 1 and the two orthogonal axes as and . The two planes and , which are perpendicular to the kagome plane, are mirror planes of the structure. They correspond respectively to K and M directions of the BZ (see inset in Fig. 1). One can see that reflection with respect to these mirror planes exchanges atoms 2 and 3 while leaving atom 1 unchanged. Therefore, the three bands created from a single orbital must combine atoms 2 and 3 in a symmetric () or antisymmetric () way to respect the mirror planes. In the first case, a contribution from atom 1 can be added (in-phase or out-of-phase), while in the second case, it must be zero. A sketch of these combinations for the different values of is given in Fig. 1(c).
We take as an example a orbital at site 1, which is even with respect to the mirror plane and odd with respect to the mirror plane. To preserve the symmetry of the kagome lattice, the orbital must rotate by 120∘ at sites 2 and 3. This is achieved by taking an appropriate combination of and orbitals [35]. In other words, we consider the even component of the / combination. The typical dispersion of the three bands obtained from a tight-binding (TB) model with only first-neighbor couplings [1] is shown in Fig. 1(b). We neglect here spin-orbit coupling (SOC) that will open a gap where two bands touch at or K [1]. This is crucial to analyse topological properties, but will not play a large role away from the crossing, where we will focus to identify the bands character. We further indicate near each dispersion the relative contributions of the three atoms in the first Brillouin zone and sketch this in Fig. 1(c). We use color to represent the band parity with respect to mirror planes perpendicular to the kagome plane and containing the high symmetry direction. The parity matches that of the orbital when sites 2 and 3 are in phase, and it is opposite when they are out-of-phase. Notably, forming a Dirac cone at K requires bands of opposite parity. However, it is often incorrectly assumed in ARPES studies of kagome compounds that band parity can be inferred directly from the dominant orbital character, leading to a wrong application of the selection rules that we describe below.
@@ III.2 CoSn calculated band structure
CoSn has a relatively simple structure with kagome planes stacked vertically on top of each other and separated by two Sn2 buffer layers [36]. This preserves the characteristic features of the simple TB model of the kagome plane, especially Dirac cones and relatively flat bands, as can be seen in the DFT calculation [14] at of Fig. 2(a). More precisely, each orbital displays a 3-band structure with very well-defined orbital character (see symbols). We label 15 the bands built from / orbitals and 710 bands from /, observable in the chosen energy window. The orbitals appear only below eV at =0 and will be ignored for the rest of the discussion (at =, a strong three-dimensional dispersion brings it much closer to the Fermi level). Two Dirac cones are formed around eV for / (bands 3 and 4) and / (bands 8 and 9). Two nearly flat bands are present near eV at for (band 7) and (band 2), with a dispersion of about 0.2 eV, much smaller than the 1 eV dispersion of the Dirac bands.
The three bands are clearly visible in this energy window, as highlighted by the marker size in Fig. 2(b). Some residual weight is observed along bands 1 and 2, which are predominantly from , evidencing hybridization between the two. The symmetry along M and K is described by the point group. This symmetry is characterized by two perpendicular mirror planes: the kagome plane itself and the perpendicular plane containing the considered symmetry direction. Since the symmetry with respect to this latter mirror plane defines the ARPES selection rules [17], we group in Fig. 2(b) the four irreducible representations of into two groups: those that are even (red) and those that are odd (blue) with respect to this mirror plane. One can check that they follow the same rules as in the TB model of Fig. 1(b).
@@ III.3 CoSn measured band structure with ARPES
We now turn to the comparison with actual ARPES experiments. In Fig. 3(a), we show a Fermi Surface for a CoSn sample lightly doped with Fe (5%), taken with linear horizontal (LH) polarization and at 126 eV photon energy, corresponding to =0 [11]. In ARPES, polarization selection rules apply when the photoelectron momentum lies in high-symmetry planes, in this case and containing either K or M depending on the sample orientation. If the polarization is even (odd) with respect to this plane, only even (odd) bands will have non-zero matrix elements [17, 18]. Fig. 3(b) sketches the experimental geometry with the orientation of LH and LV polarization in our experiment. As LH lies within the plane, it is even with respect to it, and even bands will be detected along . As LH has both even and odd components with respect to , no straightforward selection rules can be applied. On the other hand, LV polarization is along and perpendicular to , selecting even bands along and odd ones along . Note that these selection rules only predict if a band can be observed and not if its intensity will be strong.
@@ IV Unfolding the kagome lattice
This will directly impact ARPES measurements in different BZs, as it was shown for other multisite systems like graphite or graphene [20, 37] that the intensity of the ARPES bands is modulated by the sum of the TB coefficients (see also the Appendix C).
Within the TB framework, this sum can easily be calculated and is represented by marker size in Fig. 4(b-d), for each inequivalent path. One can see that only one band is strong for each of the paths, specifically the one with the three atoms in phase (see also Appendix A, which describes the same effect as a function of the hidden periodicities). This emphasizes the need to measure ARPES at least along these three paths to observe all bands. Conversely, this is a means to better identify the nature of the band observed.
More generally, when a tight-binding (TB) model is either intractable or fails to capture the relevant physics, the electronic structure can be computed at the DFT level. The resulting wave functions can then be “unfolded” [27, 39, 28] by projecting the states obtained in the true unit cell onto the reciprocal space of an alternative unit cell that accounts for additional periodicities
Such procedures are now commonly used to compare theoretical calculations with ARPES measurements when a supercell arises from an electronic ordering, such as antiferromagnetism or charge density waves. The resulting folded bands, with periodicity imposed by electronic order, typically exhibit weak intensity, in fact proportional to the strength of the perturbation [40, 41]. This approach can also be beneficial in calculations that involve large unit cells, for example, to simulate doping effects [42].
We propose here to use it to analyze the kagome lattice, where the origin of the subperiodicities is intrinsic to the multi-site structure (see Appendix A). To unfold the band structure in the large BZ (dotted line in Fig. 4(a)), we use the Fold2Bloch module [28, 43] and a transformation matrix . Figure 5 displays this calculation unfolded along the three different paths identified in Fig. 4(a). The size of the marker gives the unfolded weight corresponding to Eq. 2 and the colors refer to the orbital character, as in Fig. 2. There are well defined differences expected for spectral weights in different BZ, which will be the largest when all atoms are in-phase. For example band 1 is very strong in (a), while bands 3 and 4 are nearly zero. The opposite is true in panel (b) and (c), but differently for M and K. Generally, the unfolding effect is less clear on / orbitals, because they are more strongly hybridized with each other.
@@ V Electronic correlations in CoSn
Figure 6(a) gives an overview of most of the bands, by combining images measured under different experimental conditions in pure CoSn. Five bands (labeled 13, 7, and 8) are nearly degenerate at around 0.4 eV, and it is useful to use the combination of unfolding and parity outlined in Fig. 3 to separate them. Although SOC mixes the bands in the crossing region, obscuring the parity and orbital character definitions in this region, they remain well resolved away from for the purpose of this analysis. On the other hand, this mixing is at the origin of the nontrivial topology of CoSn, which was precisely studied in ref. 44. Our findings are in very good agreement with previous ARPES studies of CoSn [14, 11, 31], which all find a bottom position of band 1 near 0.2 eV, the flat band 7 near 0.05 eV at and the Dirac cones around 0.8 eV (black arrow). The zoom in the energy range 0.6 eV below is shown in Fig. 6(b,c) with symbols representing the experimental dispersion. Qualitatively, this agrees well with the calculated band structure, including SOC, which is shown as thin lines in Fig. 6(a). Quantitatively, however, the calculated dispersions deviate from the experimental data. The band 1 reaches minimum at 0.37 eV in calculation, which is significantly different from the experimental value of ca. 0.18 eV.
Electronic correlations are expected to renormalize the band dispersions, shrinking the effective band width [17]. We apply such a renormalization in Fig. 6(b) by dividing all dispersions by a factor of 2 with respect to . This correction improves the energy position of band 1 and its curvature, but not the other bands. Band 3 has clearly a higher binding energy than the renormalized calculation (see green arrow). Also, the Dirac cone located experimentally around 0.8 eV is now placed at 0.5 eV after the factor of 2 renormalization of the calculated band structure. In fact, we observe that the splitting between bands 1 and 3 is 0.1 eV in experiment, exactly as in the unrenormalized band dispersion with SOC. This suggests that this part of the band structure could be well described by simply shifted up the bands. We do this in Fig. 6(c), applying an upward shift of 0.18 eV. This describes the dispersions reasonably well within a larger energy window (see band 4 along M in Fig. 6(a), which is also well described by shifting up the calculation), except for band 1. Indeed, such a shift would not conserve the number of charge carriers, contrary to band renormalization that respects the Luttinger theorem [45]. As a result, the Fermi crossings of band 1, that was well described by the renormalized band structure, is not reproduced by the shifted one.
A shift is a priori not expected from electronic correlations, but could arise from surface potential [14, 12, 46]. The fact that the bulk stoichiometry is observed here, at least for at =0, first suggests that our ARPES data are bulk sensitive. Moreover, surface calculations suggest Sn surface termination for CoSn with no strong deviation from bulk [32]. We rather suggest here a novel correlation effect, where only the band crossing the Fermi level is renormalized, which makes sense as these metallic electrons should be more sensitive to the many-body effects than those of the filled bands. On the other hand, as band 1 is tied by symmetry to other bands (band 1 and band 2 are only separated by the SOC gap, for example), renormalizing band 1 necessarily raises band 2 and the bands strongly hybridized with it. It turns out that the shift corresponding to renormalizing band 1 by a factor of 2 is close to 0.18 eV, which suggests to link the two effects. Further orbital-dependent adjustments may be necessary to improve the description of bands in Fig. 6(c), but this reaches the limits of experimental accuracy. We note that a rather large range of renormalization values was given for CoSn, between 1.2 [11, 31] and 2 [14], probably depending on which features were predominantly fitted. Our hypothesis clarifies why correlations could appear as energy dependent.